Find the ordered pair which solves this system of equations: 2x-3y=11 and 4x+y=15
Answer by mananth(16946)

x Number of adult tickets
y Number of student tickets
Total attended
2 x -3 y = 11 .............1

4 x + 1 y = 15 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 3
2 x -3 y = 11
12 x + 3 y = 45
Add the two equations
14 x = 56
/ 14
x = 4
plug value of x in (1)
2 x -3 y = 11
8 -3 y = 11
-3 y = 11 -8
-3 y = 3
y = -1
(4,-1) is the solution
m.ananth@hotmail.ca